Fermi National Laboratory
 

Operation Summary for July 14, 2010 - July 16, 2010

 

Wednesday July 14

The day shift began with Operations monitoring the Tevatron (TeV) store 7974 (154.78E30) and the Recycler stash (136.3E10), with the Antiproton Source (Pbar) stacking (11.7E10), with the Linac, Booster and the Main Injector (MI) operating normally, with the MiniBooNE and Meson FTBF T-977 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, Operations monitored the accelerator complex.

During the evening shift, Operations terminated store 7974 at 4 PM. Operators prepared D0 for an access into their collision hall; they were out of access by 5:50 PM. Operations disabled T-977 at 6 PM. Operations put the TeV into HEP shot setup at 8:10 PM and established store 7977 at 9:20 PM with an average initial luminosity (AIL) of 301E30. A Booster Gradient Magnet Power Supply (GMPS) tripped off; Operations called in Booster personnel to assist in the investigation.

Thursday July 15

The midnight shift began with Operations monitoring store 7977 (175.04E30) and the stash (98.2E10), with Pbar stacking (21.7E10), and with MiniBooNE taking beam.

During the midnight shift, Operations reported that a Linac RF station (LRF5) tripped a high voltage breaker internal to its modulator. At the same time a vacuum cascade occurred in the MI-8 line. Normal beam operations resumed at 2:50 AM. At 3:40 AM, Booster extraction suffered from bad beam positions. Operators and Booster personnel determined the cause to be a problem with the pulse-forming high voltage cable for an extraction kicker (MKS04). Booster personnel took the kicker offline. Operations compensated by adjusting the remaining kicker voltages and normal beam operations resumed at 6 AM. Operations established beam to T-977 at 6:30 AM.

During the day shift, Operations terminated store 7977 at noon, put the TeV into shot setup, and established store 7978 at 1:35 PM with an AIL of 274E30. Booster personnel repaired MKS04 and brought it back online. Recycler personnel conducted proton studies. Operations disabled T-977 beam at 3:20 PM, when experimenters decided to open up MT6 section 2 for work.

During the evening shift, Operators reported problems with the MI-8 beam; both MI and Booster personnel investigated. They found evidence of transverse beam blow-up, but could not find a cause. Beam positions improved following a MI-8 line quadrupole magnet hysteresis reset.

Friday July 16

The midnight shift began with Operations monitoring store 7978 (82.93E30) and the stash (210.9E10), with Pbar stacking (21.3E10), and with MiniBooNE taking beam.

During the midnight shift, Operations terminated store 7978 at 4 AM, put the TeV into shot setup, and established store 7979 at 5:15 AM with an AIL of 264E30. Operations established beam to T-977 at 5:30 AM.

The Plans for Friday and the Weekend

The plans for today are to stack, stash, and store. The Run Coordinator will allow the following: MI studies, Pbar Debuncher scans, and Tevatron Electron Lens scans. All proton beams stop tonight at midnight in preparation for a month long shutdown, which begins on Monday, 7/19/10, at 8 AM. The last TeV store will continue colliding until midnight on Sunday.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/16/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory